OpenStudy (anonymous):

(2,3) and (4,5) write the equation of a line given 2 points on the line?

OpenStudy (lgbasallote):

lol first it was systems now it's equations of a line. your lessons are inconsistent =)) anyway.. use the two-point formula \[y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x- x_1)\]

OpenStudy (lgbasallote):

i assume you know how to use it yes?

OpenStudy (anonymous):

so its -2 and 2/2??

OpenStudy (lgbasallote):

what is?

OpenStudy (anonymous):

if you plug it in and do the solving? because 3-5 =-2 and 4/2=2? correct

OpenStudy (lgbasallote):

you went a bit inconsistent there

OpenStudy (lgbasallote):

first tell me what is the x coordinate of the first ordered pair?

OpenStudy (anonymous):

2

OpenStudy (lgbasallote):

that's your x1 what's the y coordinate of the first ordered pair?

OpenStudy (anonymous):

3

OpenStudy (lgbasallote):

that's y1 what's the x coordinate of the second ordered pair?

OpenStudy (anonymous):

4

OpenStudy (lgbasallote):

that's x2 and the y coordinate of teh second ordered pair is?

OpenStudy (anonymous):

5

OpenStudy (lgbasallote):

now you have y2 now substitute them intot he formula i wrote

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

tell me what you get later

OpenStudy (anonymous):